OpenStudy (anonymous):

Is 1/infinity = 1-0.9999.....

OpenStudy (anonymous):

anyone?

OpenStudy (experimentx):

where do you get such expression?

OpenStudy (anonymous):

i'm just think. and umm i think 0.999.. is not 1 1-0.99.. > 1-1

OpenStudy (experimentx):

0.999999999999999 and 1 are different representation of same number.

OpenStudy (experimentx):

what's on the left is quite illegal ... one does not simply divide by infinity.

OpenStudy (anonymous):

\[\frac{1}{inf}\] is less and less but it's not 0 right?

OpenStudy (anonymous):

is it kind of 0.00000..01

OpenStudy (experimentx):

\[ \lim_{x \to \infty } \frac 1x = 0\] as x keeps increasing infinity, 1/x behaves as zero ... just nut 1/infinity = 0

OpenStudy (anonymous):

but i think \[\frac{1}{\infty}=a,a \rightarrow 0\] or x*10^-n, x>0, n -> infinity

OpenStudy (anonymous):

or i have mistake?

OpenStudy (skullpatrol):

First of all infinity is not a number.

OpenStudy (anonymous):

ummm.....why

OpenStudy (anonymous):

i meant kind of "Yea, why??"

OpenStudy (skullpatrol):

A number provides particular information. Infinity is not any one particular number.

OpenStudy (anonymous):

so if infinity is what ever that infinity what is reciprocal of infinity

OpenStudy (anonymous):

cus i don't think that 1-0.99999..=0 it =0.000000....0001 and it's > 1-1 and i want to find the closest number to 1-0.99..

OpenStudy (skullpatrol):

Again just like infinity is not a number. An infinitesimal is not a number.

OpenStudy (skullpatrol):

The infinity large and the infinity small are not particular numbers.

OpenStudy (skullpatrol):

Numerals name particular numbers. They provide particular information

OpenStudy (anonymous):

and actually 0 is not number except following 0, right?

OpenStudy (skullpatrol):

What do you mean?

OpenStudy (anonymous):

can you define 0, because if a-a=nothing 0 is not value. it's variable, actually, same as infinity

OpenStudy (anonymous):

i heart from newer proof or something

OpenStudy (anonymous):

meant heard

OpenStudy (skullpatrol):

If it is 0 degrees outside then that is a number that provides particular information about the temperature just as negative degrees will provide us with a specific temperature, thus making them all numbers.